OpenStudy (anonymous):

Multiply (w+t)(w^2-wt+t^2)=

OpenStudy (anonymous):

w^3-w^2t+wt^2+tw^2-wt^2+t^3 =w^3+t^3

OpenStudy (anonymous):

Can you show me how you came up with that. I got a different answer that is why I ask.

OpenStudy (anonymous):

1st multiply with w frm 1st barckt to anather brckt ull get=w^3-w^2t+wt^2 den multiply wd t frm 1st brckt to anather brcket ull get= tw^2-wt^2+t^3 den add both of dem as ders addition in 1st brcket ull get=w^3-w^2t+wt^2+tw^2-wt^2+t^3 now -w^2t will get cancelled with +w^2t as they hav opposite signs and wt^2 will get ancelled wd -wt^2 so finally ull get =w^3+t^3 as an ans.

OpenStudy (turingtest):

when we multiply something like this out we \(distribute\) each term from the parentheses, then add them all up...